Suppose the simplified consolidated balance sheet shown below is for the entire commercial banking system. All figures are in billions of rands. The reserve ratio is 9%.


Table 2: Simplified Consolidated Banking System Balance Sheet


Assets 

Reserves    R70.5      

Securities R219.5

Loans R160


Liabilities

Deposits R450            



Based on the information above: The Money multiplier is_______. The value of excess reserves in the commercial banking system is ___ billion. If these excess reserves were loaned out, ceteris paribus, the total value of additional loans created would be___ billion. 



0.09; R30,5; R160


11,11; R30; R1778


11,11; R30; R 333


40.5; R30; R5,6


1
Expert's answer
2021-09-02T10:37:43-0400

Reserves = required reserves + excess reserves

required reserve ratio "= 9\% = 0.09"

Required reserves "= 0.09 \times 450 = 40.5"

Excess reserves "= 70.5 - 40.5 = 30"

loans created"=\frac{excess\space reserves}{reserve\space ratio}=\frac{30}{0.09}=333.33"

money multiplier"=\frac{1}{0.09}=11.11"